OpenStudy (anonymous):

A potted tree weighs 44 pounds,including the pot, the tree, and the soil.The pot, when empty weighs 2 pounds less than one third as much as the tree. The soil in the pot weighs 3 pounds less than the weight of the tree Part A: using x to represent the weight of the tree, write an expression that represents the weight of the pot. Then write an expression that represents the weight of the soil. Part B: write an equation that can be solved to find the weight of the tree.

OpenStudy (mertsj):

x+pot+soil=44 so x = 44 - pot - soil pot= (1/3)x-2 soil = x - 3

OpenStudy (anonymous):